Can you please help me this exponental problem: x^2=6x-8? Thank you!
Answer by ptaylor(2048)

Can you please help me this exponental problem: x^2=6x-8? Thank you!
x^2=6x-8
Now this is a quadratic equation and we first need to get it into standard form(Ax^2+Bx+C=0)

Subtract 6x from and add 8 to both sides of the equation:
x^2-6x+8=6x-6x-8+8 Collect like terms and we have:
x^2-6x+8=0-------------standard form where A=1, B=-6, C=8
x^2-6x+8=0 has factors (x-4)(x-2) How do I know this??? If the A coefficient of a quadratic is 1, then the B coefficient is the sum of the factors of the C coefficient, if it is factorable and the factors of 8 that work for this problem are (-4)(-2).

We now have:
x^2-6x+8=0
(x-4)(x-2)=0
x=4
x=2

Hope this helps----ptaylor